OpenStudy (anonymous):

10square root of 9x = 60

OpenStudy (anonymous):

\[10\sqrt{9x}=60\] divide each side by 10 \[\sqrt{9x}=6\] square each side \[\sqrt{9x}^{2}=6^{2} \rightarrow 9x=36\] divide both sides by 9 \[x=4\]

OpenStudy (anonymous):

Does 10square root 9x mean 10*square root 9x.Then x=4.